# Multidata Causal Discovery for Statistical Hurricane Intensity Forecasting

### v1

@@ Abstract
Improving statistical forecasts of Atlantic hurricane intensity is limited by complex nonlinear interactions and difficulty in identifying relevant predictors. Conventional methods prioritize correlation or fit, often overlooking confounding variables and limiting generalizability to unseen tropical storms. To address this, we leverage a multidata causal discovery framework with a replicated dataset based on Statistical Hurricane Intensity Prediction Scheme (SHIPS) using ERA5 meteorological reanalysis. We conduct multiple experiments to identify and select predictors causally linked to hurricane intensity changes. We train multiple linear regression models to compare causal feature selection with no selection, correlation, and random forest feature importance across five forecast lead times from 1 to 5 days (24–120 hours). Causal feature selection consistently outperforms on unseen test cases, especially for lead times shorter than 3 days. The causal features primarily include vertical shear, mid-tropospheric potential vorticity and surface moisture conditions, which are physically significant yet often underutilized in hurricane intensity predictions. Further, we build an extended predictor set (SHIPS++) by adding selected features to the standard SHIPS predictors. SHIPS++ yields increased short-term predictive skill at lead times of 24, 48, and 72 hours. Adding nonlinearity using multilayer perceptron further extends skill to longer lead times, despite our framework being purely regional and not requiring global forecast data. Operational SHIPS tests confirm that three of the six added causally discovered predictors improve forecasts, with the largest gains at longer lead times. Our results demonstrate that causal discovery improves hurricane intensity prediction and pave the way toward more empirical forecasts.
@@
: As data-driven hurricane forecasting tools become more sophisticated, it is critical to ensure that they make accurate forecasts for the right reasons, especially for tropical storms not included during training. Here, we use a technique called multidata causal discovery, which identifies a unique set of predictors driving hurricane intensity changes over the next five days by analyzing multiple realizations of the same underlying phenomenon: tropical storm in the North Atlantic. Adding these predictors to an operationally used statistical hurricane intensity prediction scheme improves forecasting accuracy. This shows that selecting predictors based on cause-and-effect, not just correlation, can result in improved forecasts that are more generalizable as well as interpretable, since the causally discovered predictors are physically consistent.
@@ 1 Introduction
Extensive research on tropical cyclones (TC), including cyclogenesis and intensification, has advanced our understanding of key atmospheric and oceanic processes gray1998formation ; elsberry1995recent ; elsberry2014advances . However the continued rise in coastal populations, and the increasing risks of wind gusts, storm surge, extreme rainfall, and severe weather pielke1998normalized ; crossett2004population underscore the need for more accurate and robust predictions of TC intensity.
Despite breakthroughs in short-range to subseasonal prediction regimes camargo2019tropical ; heming2019review ; tang2020recent , accurately predicting rapid storm intensity changes remains challenging demaria1999updated ; wang2004current ; demaria2021operational especially after 24 hours, due to error growth in initial conditions, inherent misrepresentations of the model due to lack of understanding of the underlying processes, and inadequate data assimilation techniques emanuel2016predictability . In addition to numerical weather prediction systems, statistical predictive models play a significant role in TC forecasts by contributing to the improved predictive skill of multi-model ensemble-based consensus forecasts demaria2022national . However, both statistical and dynamic models struggle to estimate the rapid intensification of TCs, which is a major challenge for storm forecasters, as multiscale air-sea interactions and radiative-convective feedback leading to rapid changes in intensity are still not fully understood wang2012recent . However, when comparing probabilistic predictions from high-resolution models and statistical schemes, the latter outperform dynamical models in predicting rapid intensification torn2021validation . The forecast skill of the Statistical Hurricane Intensity Prediction Scheme (SHIPS), which integrates large-scale predictors from climatology, persistence, and synoptic predictors to estimate hurricane intensitiesdemaria1994statistical ; MarksDeMaria2003rcliper has gradually improved over the years. SHIPS evolved from a “statistical-synoptic” to a “statistical-dynamical” model when it started using synoptic environmental conditions from dynamical models (e.g. GFS) in statistical algorithms demaria2022national . The combined consensus forecast has the best skill for TC intensity prediction, therefore SHIPS adds value to operational intensity prediction efforts cangialosi2020recent . Although SHIPS includes predictors from ocean analyses and satellite imagery, the forecast skill diminishes with lead time, partly because predictors are chosen semiempirically based on domain knowledge, climatology and persistence demaria2005further . To increase sample size, the regression coefficients for SHIPS are re-derived after each hurricane season since 1993 demaria2021operational , including annual climatological adjustments and continual updates to the predictor list. However, while the SHIPS developmental dataset provides an extensive set of environmental predictors, it may still omit key variables relevant for changes in TC intensity.
In this study, we address this challenge by adopting a causal discovery framework runge2018causal ; runge2019inferring ; runge2019detecting ; gerhardus2022causal ; runge2023causal to objectively identify and suggest new environmental predictors to be added to the latest versions of the SHIPS developmental dataset. Recent work in climate science used linear and nonlinear causal discovery methods ebert2012causal ; runge2014quantifying ; runge2019detecting to unveil spurious associations in existing statistical prediction models due to common drivers or indirect associations. Once the causal graph is known, the strength of the links between the physical variables can be determined with causal inference ricard2025causal , complementing data-driven techniques runge2023causal . In this study, we used the statistically significant causal links identified with causal discovery frameworks to determine predictors with direct causal links to changes in TC intensity.
Despite this recent progress, causality is still commonly inferred via lagged correlation analysis. The main disadvantage of lagged correlation analysis is that it does not provide insight into the causal directions of the relationships between variables mcgraw2018memory . As such, lagged correlation methods are not immune to non-causal correlations from autocorrelation effects, indirect connections through a third process, or a shared driver. Such non-causal and spurious correlations hinder interpretability in statistical prediction models Runge2014 ; barnes2015impact . Understanding the directional causality between atmospheric variables requires additional methodologies or causal inference approaches beyond the scope of lagged regression analyses. wijnands2016variable use the graphical model-based structure learning approach of the PC causal discovery algorithm to identify key variables for tropical cyclogenesis prediction. They found that predictive modeling using logistic regression employing the highest ranked variables improved statistical predictive skills; their results further suggest that causal methodologies could be used to create new tropical cyclogenesis indices. Additionally, the PCMCI+ causal algorithm — a combination of the PC algorithm and the Momentary Conditional test runge2020discovering — have been used to identify precursor regions useful for improving seasonal hurricane frequency forecasts Latos_etal2024_CausalAI .
Here, we leverage the PC algorithm and increase statistical power by applying the recently developed multidata causal discovery framework to suggest new predictors for the operational SHIPS model. This approach assumes that all Atlantic hurricane intensity changes are governed by a shared causal graph, which is reasonable given that different hurricanes represent multiple realizations of the same physical phenomenon. The multidata PC algorithm requires multivariate time series of all potential predictors for each storm, motivating our replication of the SHIPS developmental dataset using meteorological analyses, as described in Section 2. After detailing our causal methodology and experimental design in Section 3, we demonstrate in Sections 44.1 and 44.2 that causal feature selection outperforms correlation-based methods in identifying useful predictors for TC intensity forecasts, particularly at short lead times. These improvements are robust in both reanalysis-based experiments (Section 44.3), where we also show that the causal predictors nonlinearly drive intensity (Section 44.4), and in operational-like settings using real-time global model analysis and forecast fields (Section 44.6). In Section 44.5, we present a case study of Hurricane Larry to illustrate the physical relevance of the causal links between the selected predictors and short-term TC intensification.
@@ 2 Data Sources and Preprocessing
Our analysis is based on two complementary experimental designs: (i) a hierarchy of reanalysis-based experiments using the high-resolution ECMWF Reanalysis Version 5 (ERA5) hersbach2018era5 ; hersbach2020era5 , and (ii) a developmental SHIPS dataset experiment that tests the added value of these causally informed predictors alongside the operational SHIPS predictors. Our causal discovery experiments focus solely on SHIPS predictors that can be replicated using ERA5 variables; satellite-based predictors such as GOES brightness temperature and Ocean Heat Content (OHC), which are not directly reproducible in ERA5, are excluded to ensure a fair comparison and reliable interpretation of causal relationships based on ERA5 data. To define TC-centered domains and intensity change targets, we rely on the International Best Track Archive for Climate Stewardship (IBTrACS), which provides reference track and intensity data for each TC.
@@ 2.1 IBTrACS: Reference Dataset for Intensity and Track
IBTrACS knapp2010international provides a comprehensive record of TC tracks and intensity estimates compiled from all Regional Specialized Meteorological Centers and other TC warning agencies, including the Joint Typhoon Warning Center. We use IBTrACS in two complementary ways. First, the 6-hourly best-track locations define the TC-following domains used to extract area-averaged ERA5 synoptic predictors. Second, the observed maximum sustained surface (10 m) wind speeds are compared in each forecast lead time (24, 48, 72, 96 and 120 hours) to create the DELV24, DELV48, DELV72, DELV96, and DELV120 intensity change targets (in units m/s). TC intensity is commonly estimated with the satellite-based Dvorak method, which is systematically biased for weak and intense hurricanes knaff2010evaluation . To minimize potential biases related to Dvorak intensity estimates, we restrict our analysis to the North Atlantic, where routine aircraft reconnaissance missions provide regular in situ measurements of the true TC intensity. We select 247 long-lived Atlantic hurricanes (2000–2021) with lifetimes of at least four days before landfall for analysis.
@@ 2.2 ERA5 and TC PRIMED Datasets
The ERA5 predictor input to the causal feature selection algorithm consists of 6-hourly, area-averaged time series extracted from TC-following domains centered on IBTrACS best-track positions. To capture both inner-core storm dynamics and broader environmental conditions, predictors are computed for multiple radial regions around the storm center. Specifically, inner-core variables are averaged over 0–2 (approximately 0–200 km), while outer-area predictors are averaged from 200–800 km or up to 1000 km, following conventions used in the extended SHIPS developmental predictors. To supplement these ERA5-derived predictors, we also make use of the Tropical Cyclone Precipitation, Infrared, Microwave, and Environmental Dataset (TC PRIMED) Razin2023 . TC PRIMED uses ERA5 to reconstruct operational SHIPS predictors as well as environmental and thermodynamic variables across multiple pressure levels and radial regions. This dataset provides an expanded pool of predictors beyond those we directly derived from ERA5. For reference, see the documentation and the products.
@@
As the first set, we have the original operational predictors from the SHIPS developmental dataset. For causal experiments, we replicate these core SHIPS developmental predictors, which are derived directly from ERA5 reanalysis fields (e.g., VMAX, POT, PER, vertical wind shear; see Table S2 for details), providing a direct comparison to operational baselines. This is complemented by a broader pool of synoptic and thermodynamic variables sampled at multiple standard pressure levels and area-averaged within the defined radial domains, using ERA5 and supplemented by variables from TC PRIMED. These additional variables include divergence, vorticity, potential vorticity, equivalent potential temperature, geopotential height, relative humidity, air temperature, temperature gradients, precipitable water, and warm-core anomalies, among others. A complete list of variable names, radial averaging details, and pressure levels is provided in the Supplementary Information: Table S1 (original SHIPS predictors), Table S2 (ERA5 replication), Table S3 (inner-core variables), Table S4 (outer-area variables), and Table S5 (variables from TC PRIMED).
@@
A key requirement for applying causal discovery to time series is the assumption of causal stationarity, i.e., the causal relationships between variables remain invariant over time within a given TC. This allows pooling of statistical evidence across time using a sliding-window approach for conditional independence testing. Here, we also assume that the causal relationships governing TC intensity change are consistent across storms, enabling conditional independence tests to be applied across the full training set saranya2023selecting . To increase the likelihood of satisfying this assumption, we align TC time series according to their life cycle. Specifically, we aligned the Mean sea level Pressure (MSLP) time series for each storm in the training set relative to the time of minimum central pressure, smoothed using a Gaussian filter ($\sigma=3\times 6\mathrm{hr}$) to reduce noise. The idea here is to align these unique indices together to satisfy causal stationarity. For the shorter time series, we append nan values on either sides to make sure that the lengths are consistent with the longest time series with minimum pressure in the middle. We aligned the time series using these indices so that the evolution of different storms could be compared on a common reference frame. All predictors are then standardized using the mean and standard deviation computed over the training set. The final ERA5 dataset includes 214 predictors and the intensity-change target (DELV) at each forecast lead time. TCs from 2000–2019 are split into non-overlapping training and validation sets (85% and 15%, respectively), while an independent test set includes TCs from 2020–2021, along with Hurricane Wilma (2005), reserved as a particularly intense and challenging case.

### v2

@@ Abstract
Improving statistical forecasts of Tropical Cyclone (TC) intensity is limited by complex nonlinear interactions and difficulty in identifying relevant predictors. Conventional methods prioritize correlation or fit, often overlooking confounding variables and limiting generalizability to unseen TCs. To address this, we leverage a multidata causal discovery framework with a replicated dataset based on Statistical Hurricane Intensity Prediction Scheme (SHIPS) using ERA5 meteorological reanalysis. We conduct multiple experiments to identify and select predictors causally linked to TC intensity changes. We then train multiple linear regression models to compare causal feature selection with no selection, correlation, and random forest feature importance across five forecast lead times from 1 to 5 days (24–120 hours). Causal feature selection consistently outperforms on unseen test cases, especially for lead times shorter than 3 days. The causal features primarily include vertical shear, mid-tropospheric potential vorticity and surface moisture conditions, which are physically significant yet often underutilized in TC intensity predictions. We build an extended predictor set (SHIPS++) by adding selected features to the standard SHIPS predictors. SHIPS++ yields increased short-term predictive skill at lead times of 24, 48, and 72 hours. Adding nonlinearity using multilayer perceptron further extends skill to longer lead times, despite our framework being purely regional and not requiring global forecast data. Operational SHIPS tests confirm that three of the six added causally discovered predictors improve forecast skill, with the largest gains at longer lead times. Our results demonstrate that causal discovery improves TC intensity prediction and pave the way toward more empirical forecasts.
@@
: As data-driven TC forecasting tools become more sophisticated, it is critical to ensure that they make accurate forecasts for the right reasons, especially for tropical storms not included during training. Here, we use a technique called multidata causal discovery, which identifies a unique set of predictors driving TC intensity changes over the next five days by analyzing multiple realizations of the same underlying phenomenon: tropical storm in the North Atlantic. Adding these predictors to an operationally used statistical TC intensity prediction scheme improves forecasting accuracy. This shows that selecting predictors based on cause-and-effect, not just correlation, can result in improved forecasts that are more generalizable as well as interpretable, since the causally discovered predictors are physically consistent.
@@ 1 Introduction
Extensive research on tropical cyclones (TCs) has advanced our understanding of key atmospheric and oceanic processes that control cyclogenesis and intensification Gray (1998); Elsberry (1995, 2014). However, the continued rise in coastal populations and the increasing risks of wind gusts, storm surge, extreme rainfall, and severe weather Pielke Jr and Landsea (1998); Crossett et al. (2004) underscore the need for more accurate and robust TC intensity forecasts. Despite breakthroughs in short-range to sub-seasonal prediction Camargo et al. (2019); Heming et al. (2019); Tang et al. (2020), accurately predicting rapid storm intensity changes remains challenging DeMaria and Kaplan (1999); Wang and Wu (2004); DeMaria et al. (2021) especially beyond 24 hours, due to error growth in initial conditions, incomplete representation of key physical processes, and limitations in data assimilation Emanuel and Zhang (2016). For example, from 2010 - 2019, the National Hurricane Center reported mean TC intensity errors ranging from 5.7 knots ($\sim$ 2.6 m/s) for 12 hour to 12 knots ($\sim$ 6 m/s) for 48 hour forecasts Cangialosi et al. (2020). In addition to numerical weather prediction systems, statistical models play a significant role in TC forecasts by improving the skill of multi-model ensemble-based consensus forecasts DeMaria et al. (2022). Both statistical and dynamical models struggle to estimate the rapid intensification (RI) as multiscale air-sea interactions and radiative-convective feedbacks leading to rapid intensity changes are not fully understood Wang (2012). RI events account for 20 - 30% of the total intensity forecast error variance Trabing and Bell (2020), and peak intensity of rapidly intensifying TCs is routinely underestimated Torn and DeMaria (2021). Although statistical schemes outperform high-resolution dynamical models in probabilistic RI prediction Torn and DeMaria (2021), dynamical model intensity forecasts, including during RI events, improve substantially when advanced configurations such as storm-following nests are employed Alaka Jr et al. (2022).
Statistical Hurricane Intensity Prediction Scheme (SHIPS) integrates large-scale predictors from climatology, persistence, and synoptic predictors to estimate TC intensity DeMaria and Kaplan (1994); Marks and DeMaria (2003) and its forecast skill has gradually improved over time. SHIPS evolved from a “statistical-synoptic” to a “statistical-dynamical” framework through the incorporation of synoptic environmental conditions from dynamical models (e.g. GFS) DeMaria et al. (2022). The combined consensus forecast exhibits the best skill for TC intensity prediction, with SHIPS adding value to operational intensity prediction efforts Cangialosi et al. (2020). SHIPS includes an operational probabilistic RI index based on large-scale environmental predictors Kaplan et al. (2010, 2015). Machine learning approaches have further improved 24 hour RI prediction skill, with SHIPS-compatible extensions using satellite-derived predictors, hybrid deep learning–gradient boosting frameworks, and operational consensus models Su et al. (2020); Boussioux et al. (2022); Ko et al. (2023).
Although SHIPS incorporates predictors from ocean analyses and satellite imagery, forecast skill decreases with lead time, partly because predictors are chosen semi-empirically based on domain knowledge, climatology, and persistence DeMaria et al. (2005). To increase sample size, SHIPS regression coefficients are re-derived annually with climatological adjustments and updates to the predictor list DeMaria et al. (2021). While the SHIPS developmental dataset provides an extensive set of environmental predictors, it may still omit key variables relevant for changes in TC intensity.
In this study, we address this limitation by adopting a causal discovery framework Runge (2018); Runge et al. (2019a, b); Gerhardus and Runge (2022); Runge et al. (2023) to objectively identify new environmental predictors for inclusion in the operational SHIPS developmental dataset. In causal discovery, relationships among variables are represented using a causal graph, where nodes correspond to physical variables, and directed edges denote direct causal influences, potentially with time lags. While causal discovery aims to infer the full causal structure among a set of variables, causal feature selection instead uses a causal discovery algorithm to identify a subset of physically meaningful predictors that are most relevant for predicting a target variable, without reconstructing the full causal graph. Climate science studies have applied linear and nonlinear causal discovery methods Ebert-Uphoff and Deng (2012); Runge et al. (2014b, 2019b) to identify spurious associations in statistical prediction models arising from common drivers or indirect associations. Once the causal graph is known, the strength of the links between the physical variables can be determined with causal inference Ricard et al. (2025), complementing data-driven techniques Runge et al. (2023). Despite these advances, causality is still commonly inferred using lagged correlation analysis, which does not resolve the directionality of the relationships between variables McGraw and Barnes (2018). As such, lagged correlation methods are susceptible to non-causal correlations from autocorrelation effects, indirect connections through a third process, or a shared driver hindering interpretability in statistical prediction models Runge et al. (2014a); Barnes and Screen (2015). Understanding the directional causality between atmospheric variables hence requires additional methodologies or causal inference approaches beyond the scope of lagged correlation analyses. Several studies have demonstrated the potential of causal discovery methods in tropical meteorology. For example, Wijnands et al. (2016) used the graphical model-based structure learning approach of the Peter-Clark (PC) causal discovery algorithm to identify key variables for tropical cyclogenesis prediction and found that predictive modeling using logistic regression employing the highest ranked variables improved statistical predictive skills. Additionally, the PCMCI+ causal algorithm — which combines the PC algorithm with Momentary Conditional Independence test Runge (2020) — has been used to identify precursor regions useful for improving seasonal TC frequency forecasts Latos et al. (2024). Motivated by these advances, we apply the PC algorithm within a multidata causal discovery framework to identify statistically significant causal links and select predictors with direct causal influence on TC intensity change for the operational SHIPS model. This approach treats individual Atlantic tropical cyclones as multiple realizations of the same underlying physical process, assuming that their intensity evolution shares a common causal structure that can be represented by a single causal graph. The multidata PC algorithm requires multivariate time series of all candidate predictors for each storm, motivating our replication of the SHIPS developmental dataset using meteorological analyses, as described in Section 2. After detailing the causal methodology and experimental design in Section 3, we demonstrate in Sections 44.1 and 44.2 that causal feature selection outperforms other feature selection methods, particularly in short-range time-scales. These improvements are robust in both reanalysis-based experiments (Section 44.3), where we also show that the causal predictors nonlinearly drive intensity (Section 44.4), and in operational-like settings using real-time global model analysis and forecast fields (Section 44.6). In Section 44.5, we present a case study of Hurricane Larry to illustrate the physical relevance of the causal links between the selected predictors and short-term TC intensification.
@@ 2 Data Sources and Preprocessing
Our analysis is based on two complementary experimental designs: (i) reanalysis-based experiments using the high-resolution ECMWF Reanalysis Version 5 (ERA5) Hersbach et al. (2018, 2020), and (ii) experiments using SHIPS developmental to test the added value of causal predictors compared to the operational SHIPS predictors. To define TC-centered domains and intensity change targets, we rely on the International Best Track Archive for Climate Stewardship (IBTrACS), which provides reference track and intensity data for each TC.
@@ 2.1 IBTrACS: Reference Dataset for Intensity and Track
IBTrACS Knapp et al. (2010) provides six-hourly best-track positions and intensity estimates. Best-track locations define TC-following domains used to extract area-averaged ERA5 predictors, while maximum sustained surface (10 m) wind speeds are used to compute intensity change targets at forecast lead times of 24, 48, 72, 96 and 120 hours (DELV24, DELV48, …, DELV120 in m/s). To minimize potential biases related to Dvorak intensity estimates, we restrict our analysis to the North Atlantic basin, where routine aircraft reconnaissance missions provide regular in situ measurements of TC intensity. We select 247 long-lived Atlantic TCs from 2000 to 2021 with lifetime of at least four days prior to landfall.
@@ 2.2 ERA5 and TC PRIMED Datasets
Causal discovery is restricted to SHIPS predictors that can be replicated using ERA5; satellite-only predictors such as GOES brightness temperature and ocean heat content are excluded to ensure consistency. The ERA5 predictor inputs consist of six-hourly, area-averaged time series extracted from TC-following domains centered on IBTrACS best-track positions. Predictors are computed for multiple radial regions around the storm center to capture inner core and outer area conditions: inner-core variables are averaged over 0–2 (approximately 0–200 km), while outer-area predictors are averaged from 200–800 km or up to 1000 km, following conventions used in the extended SHIPS developmental dataset. To avoid backward-in-time causal dependencies, we only consider predictors at analysis (00 hour) time unlike operational SHIPS setup, which incorporates dynamical model forecast fields. As a result, forecast skill at longer lead times may be underestimated when synoptic-scale evolution is important. To expand the predictor set further, we use the Tropical Cyclone Precipitation, Infrared, Microwave, and Environmental Dataset (TC PRIMED) Razin and others (2023). TC PRIMED uses ERA5 to reconstruct operational SHIPS predictors as well as environmental and thermodynamic variables across multiple pressure levels and radial regions. An important limitation of the area-averaging approach is that it does not capture azimuthal asymmetries, which may be relevant for sheared or rapidly intensifying TCs Hazelton et al. (2017); Rogers et al. (2013); McNeely et al. (2020) and could be considered in future experiments.
@@
As the first set, we have the original operational predictors from the SHIPS developmental dataset (Table S1). For causal experiments, the predictor set includes replicated SHIPS variables derived from ERA5 reanalysis fields (e.g., VMAX, POT, PER, vertical wind shear; see Table S2 for details), directly comparable to operational baselines. This is complemented by a broader set of synoptic and thermodynamic variables sampled at multiple pressure levels and radial domains. These additional variables include divergence, vorticity, potential vorticity, equivalent potential temperature, geopotential height, relative humidity, air temperature, temperature gradients, precipitable water, and warm-core anomalies, among others. Complete variable definitions, radial averaging details, and pressure levels are provided in Tables S1–S5 of the Supplementary Information.
@@
A key requirement for applying causal discovery to time series is the assumption of causal stationarity, i.e., the causal relationships between variables remain invariant over time within a given TC and across multiple storms in the multi-data setting. This allows pooling of statistical evidence across time and the full training set for conditional independence testing Ganesh et al. (2023). Accordingly, TC time series are aligned relative to the time of minimum central pressure to satisfy the causal stationarity assumption. The mean sea level pressure (MSLP) time series for each storm in the training set are smoothed using a Gaussian filter ($\sigma=3\times 6\mathrm{hr}$) to reduce noise and the time of minimum central pressure is used as the reference point for alignment. For the shorter time series, we append NaN values on either sides to make sure that the lengths are consistent with the longest time series with minimum pressure in the middle so that the evolution of different storms could be compared on a common reference frame. All predictors are subsequently standardized using the mean and standard deviation computed from the training set. The final ERA5 dataset includes 214 predictors and the intensity-change target (DELV) at each forecast lead time. A total of 216 TCs from 2000–2019 are used within a storm-based cross-validation framework with predefined, non-overlapping training and validation splits, while an independent test set comprises 31 TCs from 2020–2021 including Hurricane Wilma (2005), retained as a particularly intense and challenging case.